OpenStudy (anonymous):

In a right triangle, if one of the angles is 35°, what are the measures of the other two angles?

OpenStudy (anonymous):

one is 90 one is 35 3rd?

pooja195 (pooja195):

Hi Welcome to openstudy :) In a right triangle you will always have a 90 degrees angle since there is a right triangle present call the other missing angle x the 3 angles must add to 180 degrees, so.... m<1 + m<2 + m<3 = 180 \[\huge~\rm~\bf~35+90+x = 180\]solve for x

OpenStudy (alexandervonhumboldt2):

 \[\Huge\mathbb {{x+35+90=180}}\] solve for x. \[\Huge\mathbb {{x=180-35-90=?}}\]